A triangle has three angles, R, S, and T. Angle T is 40° greater than angle S. Angle R is 8 times angle S. What is the measure of each angle?
Answer by venugopalramana(3286)

A triangle has three angles, R, S, and T. Angle T is 40° greater than angle S.
T=S+40
Angle R is 8 times angle S.
R=8S
What is the measure of each angle?
IN A TRIANGLE SUM OF 3 ANGLES =180
R+S+T=180
8S+S+S+40=180
10S=180-40=140
S=140/10=14
T=14+40=54
R=8S=8*14=112